you have heard about three different vectors to clone DNA fragments (plasmids, BACs, and YACs). Which of these would be most appropriate to clone a DNA fragment with the length of 5,000 bp, and which would be most appropriate for a fragment of 200,000 bp?

The genome of fruit flies has a size of about 175 Mbp. Assuming that each base pair has a height of 3 Angstrom (1 Angstrom is 10^-10 m), how long is the genome of a fruit fly in meters / millimeters?
Answer #1

Q.1

Answer:

  • 5,000 bp = Plasmid
  • 200,000 = BAC / YAC

Explanation:

Maximum insert size limit:

Plasmid = 10kb (10,000 bp)

BAC = 300 Kb (300,000 bp)

YAC = 1400 Kb (1,400,000 bp)

------------------------------------------------------------------------------------------------------------------------

Q.2

Solution:

175 Mbp = 175,000 Kbp

175,000 Kbp = 175,000,000 bp

> 175,000,000 * 3 = 525,000,000 Angstrom

> 525,000,000 * 10^-10 m = 0.0525 m

Answer: 0.0525 meter